OpenStudy (anonymous):

A square of perimeter 20 is inscribed in a square of perimeter 28. What is the greatest distance between a vertex of the inner square and a vertex of the outer square?

OpenStudy (lgbasallote):

im guessing \(\sqrt{49 + 49/4}\) but i guess im wrong :/

OpenStudy (anonymous):

\[5\sqrt{2}+1/2 (7\sqrt{2}-5\sqrt{2})\]

OpenStudy (anonymous):